A certain affects virus 0.9% of the population. A test used to detect the virus in a person is positive 87% of the time if the person has the virus (true positive) and 10% of the time if the person does not have the virus (false postive). Fill out the remainder of the following table and use it to answer the two questions below.

Infected Not Infected Total
Positive Test
Negative Test
Total 900 99,100 100,000

a) Find the probability that a person has the virus given that they have tested positive. Round your answer to the nearest tenth of a percent and do not include a percent sign.
P(Infected | Positive Test)= %
b) Find the probability that a person does not have the virus given that they test negative. Round your answer to the nearest tenth of a percent and do not include a percent sign.
P(Not Infected | Negative Test) = %
Answer #1

a.

P(a person has the virus given that they have tested positive) = 783/10693 = 0.07322 = 7.3%

b.

P(Not Infected | Negative Test) = 89190/89307 = 0.9986 = 99.87%